At the end of the year, overhead applied was $3,706,000. Actual overhead was $3,418,000. Closing over/underapplied overhead into Cost of Goods Sold would cause net income to

a.decrease by $576,000

b.decrease by $288,000

c.increase by $288,000

d.increase by $576,000

A manufacturing company applies factory overhead based on direct labor hours. At the beginning of the year, it estimated that factory overhead costs would be $345,800 and direct labor hours would be 42,600. Actual manufacturing overhead costs incurred were $320,800, and actual direct labor hours were 55,000. What is the predetermined overhead rate per direct labor hour?

a.$12.18

b.$6.49

c.$9.74

d.$8.12

In a process cost system, the cost of completed production in Department A is transferred to Department B by which of the following entries?

a.debit Finished Goods—Dept. A; credit Work in Process—Dept. B

b.debit Work in Process—Dept. B; credit Cost of Goods Sold—Dept. A

c.debit Work in Process—Dept. B; credit Work in Process—Dept. A

d.debit Work in Process—Dept. B; credit Finished Goods—Dept. A

A manufacturing company applies factory overhead based on direct labor hours. At the beginning of the year, it estimated that factory overhead costs would be $327,544 and direct labor hours would be 40,943. Actual factory overhead costs incurred were $395,151, and actual direct labor hours were 51,452. What is the amount of overapplied or underapplied manufacturing overhead at the end of the year?

a.$411,616 overapplied

b.$16,465 underapplied

c.$16,465 overapplied

d.$84,072 underapplied
Answer #1
overhead applied-actual overhead
3706000-3418000 = 288000 increase
Option "d" is CORRECT
Predetermine overhead rate = estimated factory overhead/direc labor
predetermine overhead rate = 345800/42600 = 8.12
Option "D" Is CORRECT
Journal entry would be
Particulars Debit Credit
work in process-Dept B XXX
work in process-Dept A XXX
Option b is Correct
overhead rate per Direct labor hr = estimated factory overhead/estimated direct labor hr
overhead rate per direct labor hr = 327544/40943 = 8
actual direct labor hr = 51452
applied factory overhead = 51452*8 = 411616
actual factory overhead = 395151
overapplied overhead = (411616-395151) = 16465 overapplied
Option "c" is CORRECT

Please give a thumbs up if it is helpful & let me know if any doubt